OpenStudy (anonymous):

Conversion: 72000 cm/hr^2 = ? m/s^2 Conversion: 500 km/minute = ? m/s

OpenStudy (anonymous):

500 km/min=(500*1000 m)/60 sec 500 km/min=500000m/60s 500 km/min=(500000/60)m/s oooooooooooooooooooooooooooooooo